If at the moment t=0 the sepration between the car is 100 m find the sepration between the car at t=2 s.. Given velocity of one with respect to other is 25 m /s

Grade:12th pass

2 Answers

Vikas TU
6 years ago
Dear Student,
At t = 0 , partition between auto = 100 m 
At t = 2 sec speed is 25 m/sec 
Separation = speed * time 
= 25 * 2 
= 50 m .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
navaneeth praneeth
2 years ago
At time t=0 seperation is 100m 
》their of each car is 50/2 =25m/s
》separation distance=speed×time
》25×2=50m